OpenStudy (anonymous):

An arc of 6.5cm subtends a central angle of 45 degrees. Find the length of the radius of the circle.

OpenStudy (anonymous):

\[2 \pi r \frac{45}{360}=6.5\]

OpenStudy (anonymous):

\[28/2\pi\]